Session: Mar 2008 (session 1)
Wavelength: 3.6 cm
To give feedback on this experiment, click
Medicina (Roda at Sun Mar 9 07:11:44 2008) Success :
Onsala (Michael Lindqvist at Sun Mar 9 09:51:58 2008) Minor failures :
* First scan recorded on a disk going to JIVE.
* Partly bad weather (rain)
* Some Mark5 problems (day 68, morning) of the type:
2008.068.10:07:11.40?ERROR m5 -900 Previous recording still active
2008.068.10:07:11.41?ERROR m5 -906 MARK5 return code 6: inconsistent or conflicting request
2008.068.10:48:43.37?ERROR m5 -900 Previous recording still active
2008.068.10:48:43.37?ERROR m5 -906 MARK5 return code 6: inconsistent or conflicting request
2008.068.10:53:01.36?ERROR m5 -900 Previous recording still active
see log for details. Disk pack changed.
Hartebeesthoek (Jonathan Quick at Mon Mar 10 08:03:48 2008) Success :
No known problems, some light rain at one stage.
Robledo (Cristina Garcia-Miro at Mon Mar 10 08:38:29 2008) Success :
Minor S-band RFI at VC01 (schedule lines #75, #262, #1061) and VC03 (schedule lines #228, #466).
Effelsberg (Uwe Bach at Mon Mar 10 11:14:13 2008) Minor failures : LATE
Started at 13:10 UT, because of snow in the morning.
Noto (P. Cassaro at Thu Mar 13 12:43:40 2008) Minor failures : RECE LATE
The observation has started normally, but the cal diode was out of order, so no Tsys informations are available, until 10:00 UT. Once fixed the Tcal, the bbc #8 has been unlocked for the rest of the observation.